OpenStudy (anonymous):

What is the volume of a cube whose sides each have a measure of 10 cm? 600 cm3 800 cm3 900 cm3 1000 cm3

OpenStudy (anonymous):

uses the cube formula which is v=side^3

OpenStudy (anonymous):

\[10^{3}\]

OpenStudy (anonymous):

oh I get it thanks!

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

volume of a cube=a*a*a=10*10*10=1000 cubic centimeters